OpenStudy (unklerhaukus):

Solve the second order DE [x^2y^{\prime\prime}+3xy^\prime=2x\]

OpenStudy (unklerhaukus):

\[x^2y^{\prime\prime}+3xy^\prime=2x\] \[\text {let}\quad \ln x=t \]\[x=e^t\]\[\frac{\text d x}{\text dt}=e^t\] \[e^{2t}y^{\prime\prime}+3e^ty^\prime=2e^t\]\[y^{\prime\prime}+(3-1)y^\prime=2e^t\]\[y^{\prime\prime}+2y^\prime=2e^t\]

OpenStudy (unklerhaukus):

\[y^{\prime\prime}_c+2y_c^\prime=0\] \[m^2+2m=0\]\[m(m+2)=0\]\[m=0,-2\] \[y_c=A+Be^{-2t}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+2y_p^\prime=2e^t\] \[y_p=Ce^t;\qquad y_p^\prime=Ce^t;\qquad y_p^{\prime\prime}=Ce^t\] \[Ce^t+2Ce^t=2e^t\]\[3Ce^t=2e^t\]\[C=\frac 23\]\[y_p=\frac23e^t\]

OpenStudy (unklerhaukus):

\[y(t)=A+Be^{-2t}+\frac 23e^t\] \[y(x)=A+Bx^{-2}+\frac 23x\]

OpenStudy (unklerhaukus):

\[\Large\color\red\checkmark\]